A firefighting crew uses a water cannon that shoots water at 28.0 m/s at a fixed angle of 47.0 ∘ above the horizontal. The firefighters want to direct the water at a blaze that is 12.0 m above ground level. How far from the building should they position their cannon? find D1 and D2
Answer #1

given that

initial speed vi = 28 m/s

angle = theta = 47 degrees

height y = 12 m

from the relation

y = voy*t-1/2*gt^2

12 = 28*sin47*t-1/2*9.8*t^2

12 = 20.48t-4.9t^2

4.9t^2-20.48t+12 = 0

it is a quadratic equation like ax^2+bx+c = 0

roots are

t = 0.705 s and 3.475 s

D1 = vox*t1 = 28*cos(47)*0.705 = 13.46 m

D2 = 28*cos(47)*3.475 = 66.4 m